OpenStudy (anonymous):

MEDAL REWARD :) At the city museum, child admission is $5.80 and adult admission is $9.60 . On Monday, twice as many adult tickets as child tickets were sold, for a total sales of $675.00 . How many child tickets were sold that day?

OpenStudy (nurali):

Let's have x be the number of children tickets and 2x will be the adults. Now we can set up the following equation: 5.8x+9.6(2x)=675 Expand this out: 5.8x+19.2x=675 Combine like terms: 25x=675 Divide both sides by 25 to get: x=27 There we go, 27 children tickets were sold.

OpenStudy (anonymous):

awesome thank you !!

OpenStudy (anonymous):

What about this one? A Web music store offers two versions of a popular song. The size of the standard version is 2.9 megabytes (MB). The size of the high-quality version is 4.7 MB. Yesterday, there were 1260 downloads of the song, for a total download size of 4500 MB. How many downloads of the high-quality version were there?

OpenStudy (nurali):

Let x= number of standard versions downloaded 1260-x = number of HQ versions downloaded 2.9x+4.7(1260-x)=4500( in MB ) 2.9x+5922-4.7x+=4500 2.8x=5922-4500 2.8x=1422 x=508 1260-508=752 There were 752 downloads of the HQ version